Review of ⟨Title of Presentation⟩ Presented by ⟨Name of Speaker⟩ to CS501 ⟨Date of Presentation⟩

⟨Name⟩

Department of Computer Science and Engineering State University of New York at Buffalo 226 Bell Hall Buffalo, NY 14260-2000

                           ⟨login⟩@cs.buffalo.edu

⟨date of review⟩

1 Summary

Give a short summary of the presentation in your own words, to the best of your ability.

2 Reaction

Briefly describe your reaction to/impression of the presentation. How interesting do you find the research area as a potential area for your own future work?

Anonymous Feedback to

⟨Name of Speaker⟩ on your Presentation to CS501 of

⟨Date of Presentation⟩

Critique of Contents

Critique of Style